The solubility of O2 in water is approximately 0.00380 g L-1 when the temperature is 25.0...

  1. The solubility of O2 in water is approximately 0.00380 g L-1 when the temperature is 25.0 and the partial pressure of gaseous oxygen is 1.00 atm. What will the solubility of oxygen be if the oxygen pressure is readjusted to 1.00 x 103 torr?
Answer #1

Solubility=KP

K is called henrys constant

P is the partial pressure

At constant temperature solubility is directly proportional to the partial pressure

S1/S2=P1/P2

P1= 1 atm

P2= 1000 torr= (1000/760) =1.31578947368 atm

S1=0.00380 g/L

S2=(S1*P2) /P1

S2=(0.00380 g.L-1*1.315789 atm) /1 atm

S2=0.00499 g. L-1